The model y = 100 + 2x1 - 4x2 + 8x3 is used to predict a company's sales (y) in a given quarter. x1, x2, and x3 are binary variables that represent quarter 1, quarter 2, and quarter 3, respectively. What is the predicted level of sales for quarter 2?

108, 100, 96, or 102
Answer #1

Given model is,

y = 100 + 2x1 - 4x2 + 8x3

x1, x2 and X3 are binary variables that represent quarter 1, quarter 2 and quarter 3 respectively.

We want to predict the level of sales for quarter 2.

y = 100 + (2 * 0) - (4 * 1) + (8 * 0)

=> y = 100 + 0 - 4 + 0

=> y = 96

Therefore, the predicted level of sales for quarter 2 is 96